Derive the following relation for flow through variable area duct:

( ) 2 = โˆ’1 dA dV M

A V

Where A is cross sectional area, V is velocity of flow and M is Mach number. Also discuss the

significance of this equation.


1
Expert's answer
2021-03-15T11:18:10-0400

The area velocity realtionship for incompressible fluid is given by the continuity equation as

"A\times V = Constant" .

From the above equation it is clear that with the increase of area, velocity decreases. But in case of compressible fluid equation is given by


"\rho\times AV = Constant"

Differentiating above equation we get,


"\rho [AdV + VdA] + AVd\rho = 0"

or

"\rho AdV + \rho VdA + AVd\rho = 0"


Dividing by "\rho AV" , we get


"\dfrac{dV}{V} + \dfrac{dA}{A} + \dfrac{d\rho}{\rho} = 0"

The Eulers equation for compressible fluid is given by,


"\dfrac{dp}{\rho} + VdV + gdz = 0"

Neglecting the Z term, Hence we can write the equation.

"\dfrac{dp}{\rho}\times \dfrac{d\rho}{d\rho} + VdV = 0"


"\dfrac{dp}{d\rho} = C^2"

Hence the above equation becomes as


"C^2\dfrac{d\rho}{\rho} + VdV = 0"

"\dfrac{d\rho}{\rho} = -\dfrac{VdV}{C^2}"

Substituting the value of "\dfrac{d\rho}{\rho}" in equation, we get


"\dfrac{dV}{V} + \dfrac{dV}{A} - \dfrac{VdV}{C^2} = 0"


"\dfrac{dA}{A} = \dfrac{VdV}{C^2} - \dfrac{dV}{V}"


"\dfrac{dA}{A} = \dfrac{dV}{V}[M^2 - 1]"